# 1CIBM – Center for Biomedical Imaging, Switzerland

### v1

@@ Abstract
Quality control (QC) has long been considered essential to guarantee the reliability of neuroimaging studies. It is particularly important for fetal brain MRI, where acquisitions and image processing techniques are less standardized than in adult imaging. In this work, we focus on automated quality control of super-resolution reconstruction (SRR) volumes of fetal brain MRI, an important processing step where multiple stacks of thick 2D slices are registered together and combined to build a single, isotropic and artifact-free T2 weighted volume. We propose , a machine-learning method that extracts more than 100 image quality metrics to predict image quality scores using a random forest model. This approach is well suited to a problem that is high dimensional, with highly heterogeneous data and small datasets. We validate in an out-of-domain (OOD) setting and report high performance (ROC AUC = 0.89), even when faced with data from an unknown site or SRR method. We also investigate failure cases and show that they occur in of the images due to ambiguous configurations for which the rating from the expert is arguable. These results are encouraging and illustrate how a non deep learning-based method like is well suited to this multifaceted problem. Our tool, along with all the code used to generate, train and evaluate the model will be released upon acceptance of the paper.
@@
Although the quality of SRR volumes greatly influences subsequent processing steps like automatic tissue segmentation and surface extraction, only a few methods have been proposed, and their validation was limited. Closest to this work, the study of [18] used multi-instance deep learning methods for automated quality control on 3D volumes reconstructed using only the approach from [10]. The model trained in this work is not publicly available and the validation was restricted to one acquisition site. Other works have taken a more descriptive approach to the quality of SRR volumes. The work of [19] focused on one popular SRR method called NiftyMIC [14], and proposed to predict the quality of the reconstructed volume based on the percentage of slices from the input stacks that NiftyMIC uses for reconstructing that volume. The work of [17] proposed a protocol for fetal brain MRI quality assessment based on the evaluation of several image and artifact properties, alongside an overall image quality score. However, their sample size remained quite small and they did not perform an automated quality assessment. Finally, while a few automated QC methods have been proposed for low-resolution fetal brain MRI [20, 21, 22], they do not readily apply to SRR volumes, since the artifacts can be quite different.
In this work, we propose to tackle the challenge of automated QC for multi-SRR, multi-site data. Our contributions are as follows: (i) We collect a unique dataset consisting of 673 manual quality ratings combining multiple acquisition sites and multiple SRR methods (ii) We build upon FetMRQC [22] to propose a robust approach based on random forest. We show that this non deep-learning approach, when combined with appropriate data pre-processing, provides a high performance, while requiring only a modest amount of training data. (iii) We evaluate the out-of-domain robustness of our method, focusing on unseen sites and SRR methods. (iv) We study the failure modes of our model and show that 45% of the failures are due to ambiguous ratings, and that nearly all false negatives (wrongly flagged as bad quality) have some level of artifact that, although it is not disturbing to a human rater, is picked up by . The code will be made available upon acceptance of the work. We believe that this method will serve as first open source, easy-to-use baseline for future works in the field.
@@ 2.1 Demographics
Data. Fetal brain MRI examinations were retrospectively collected from ongoing research studies at three hospitals: H1 (Lausanne University Hospital, n=141), H2 (BCNatal, n=180) and H3 (University Children’s Hospital Zürich (KISPI), n=37). The study received ethical approval from each center’s institutional review board. These 438 subjects were acquired on different Siemens scanners (Erlangen, Germany) at 1.5T or 3T for H1 and H2, and on different GE Healthcare (Chicago, USA) scanners for H3. As the aim of quality control is to be applicable to a wide range of scenarios, each subject was reconstructed using a randomly assigned set of parameters: a randomized SRR method, target resolution () and selection of input stacks. This resulted in 673 reconstructed volumes using four widely-used SRR methods: SVRTK [11, 23] (n=238), NeSVoR [15] (n=213), NiftyMIC [14] (n=202), MIALSRTK [13] (n=20).
Manual quality ratings. After the SR reconstructions were completed, a manual quality assessment was performed following the protocol of Sanchez et al. [17] (we refer readers to that paper for details on the protocol). The protocol was slightly adapted to focus on categories that were found to have an inter-rater reliability of , leading to the categories illustrated on the right of Figure 1. Global quality rating was obtained as a continuous variable between 0 and 4, and the exclusion threshold was set to 1. Two experts, R1 (20 years of experience in medical image analysis) and R2 (4 years of experience) performed manual quality annotations on all the reconstructions (323 for R1 and 448 for R2, with 98 cases rated twice). The annotation interface allowed to streamline the process, and ratings took a total of 11h40 (7h10 for R1, 4h30 for R2), with a median time of 47 seconds per case. Details on the distribution of SRR quality for the three main methods are provided in Figure 2.
@@ 2.2 Automated quality control
Image quality metrics. IQMs are features capturing complementary statistics on a given volume. Well-designed IQMs correlate with various image artifacts, enabling the training of a predictive QC model. However, IQMs that are commonly used in adult MRI are not readily applicable to fetal SRR volumes, as they can rely for instance on air as a background [5], which is inappropriate in utero as the fetus is surrounded by maternal tissues. This prompted us to adapt the IQMs of FetMRQC [22] designed for stacks of 2D slices of T2w fetal brain MRI to the SRR setting. These IQMs are based on image intensity (e.g. mean intensity, SSIM of neighbouring slices, etc.), mask statistics (brain mask centroid coordinates) or segmentation-related statistics (tissue-wise statistics like volume, or contrast between tissue intensities). In addition to these metrics we incorporate new topological IQMs based on Betti numbers and Euler characteristics [24], as this has been shown to be an important metric to quantify the quality of a fetal brain MRI segmentation [25]. Since several IQMs required a brain mask and tissue segmentation, we used the state-of-the-art BOUNTI [26] to compute these maps. In total, 106 IQMs were computed (82 from FetMRQC and 24 from topology). We will provide a detailed description of the IQMs used along with the code.
@@
Predictive model. is based on a random forest, and trained in a binary classification setting, where it predicts whether a data sample has quality below the threshold of , and should be excluded. In addition to it, we considered two variants of . The first one was trained by re-weighted samples according to their underlying continuous label: samples with a quality rating close to 0 (bad quality) or 4 (excellent quality) were given weights of 1, while samples close to the decision threshold (quality score = 1) were given a weight of 0. It is referred to as reweighting. The second one, called Predict_artifacts, made use of the available manual ratings of the type of artifacts present: based on the IQMs, a random forest regressor was trained to predict each specific artifact (cf. Figure 1 right), and then these predictions were used as additional IQMs for the prediction.
@@ 2.3 Experiment design
2.3.1 Evaluation setting
Previous studies have shown that domain shift is a major obstacle to the generalization of ML-based methods in medical imaging [27], motivating us to evaluate our models in four different experiments. In-domain (ID), where the data is split by subject across centers; leave-one-site-out cross-validation (CV), where all the data from a single site were used for testing; leave-one-SRR-out cross-validation, where all the data from a given SRR were used for testing; leave-one-site/SRR-out, where all the data from a given site and a given SRR method were left out during training, and the model was tested on the combined data of the left-out site and SRR. This simulates a complex domain shift, where our method is tested on a new site, with data reconstructed using an unseen SRR method. Each experiment was repeated 10 times to account for random variations in the model training and training data sampling.
@@ 3 Results
We first carried out an in-domain ablation study on the various preprocessing options and variants of . We also considered various thresholds for exclusion from the model’s predicted probabilities, and investigated how changing impacted BA, sensitivity and specificity. Results are presented in Table 1, where we see that using segmentation and IQMs pre-processing reaches the best performance with AUC=0.93 and BA=0.85. Choosing increases the specificity of the model, and we chose to keep it at that level for the rest of the paper. The addition of topological metrics improved the AUC and BA by 3% and 4% respectively. Finally, when looking at the oracle, which uses manual ratings of artifacts as the basis for prediction, we see that our models come close to the predictive ability of the oracle, but fall short by a margin of approximately 7-8%. We also experimented with the site-SRR domain shift removal using ComBaT [28], robust scaling, alternative classifiers like logistic regression, and nested cross-validation but these approaches did not improve the performance and are not reported.
@@
In this experiment, we evaluate our model on three OOD settings: leave-one-site-out, -SRR-out and -site/SRR-out. Our results, in Table 2, show an increasing drop in performance across settings: it is less challenging to generalize to an unseen site than to an unseen SRR, and the most challenging setting is the double shift site+SRR. Although the AUC and BA do not vary much for with robust preprocessing, there is a large drop in sensitivity when evaluating the model on a different site (-7%) and SRR (-11%) and a drop in both sensitivity and specificity in the site+SRR case (resp. -9%; -13%). The Site+SRR case also shows a larger variance in sensitivity and specificity of models, suggesting that the model might exhibit some degree of unreliability on this challenging domain shift. Nevertheless, generally retains the best average performance.
@@
Out of 673 cases, the prediction was consistently false for 95 (14%)across the repeated experiments. Figure 3 breaks down the distribution of the errors as a function of the manual quality rating. Out of the 95 failed cases, 23 correspond to cases with a strong bias field, and 22 to a low tissue contrast (Figure 3D). Both were challenging for ’s IQMs to capture and accurately classify. This is one limitation of feature engineering approaches that cannot capture all artifacts encountered. This is further illustrated by the cases with streaks and cut brain (Figure 3A and B), that are confidently predicted as good quality, while being excluded by raters. Nevertheless, out of 95 failed cases, 42 had reference quality scores close to the decision boundary and we consider them as ambiguous cases, where a different rater could have assigned a different label (cf. Figure 3C). These cases might be interpreted as limitations from the ground truth data as well as a challenging cases for . Lastly, systematic failure cases also included images of obviously pathological cases, for which the automated segmentation method failed, leading in turn to inaccurate IQMs and mis-classification by the model.
@@ 4.0.1 Acknowledgements
This work was funded by Era-net NEURON MULTIFACT project (TS: Swiss National Science Foundation grant 31NE30_203977; AM, GA: French National Research Agency, Grant ANR-21-NEU2-0005; EE: Instituto de Salud Carlos III (ISCIII) grant AC21_2/00016 , GMJ: Ministry of Science, Innovation and Universities: MCIN/AEI/10.13039/501100011033/), the SNSF project 215641, and the SulcalGRIDS Project, (GA: French National Research Agency Grant ANR-19-CE45-0014). AJ is supported by the Prof. Max Cloetta Foundation, EMDO Foundation and Vontobel Foundation. TS, VZ and MBC acknowledge access to the facilities and expertise of the CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by CHUV, UNIL, EPFL, UNIGE and HUG.

### v4

@@ Abstract
Quality control (QC) has long been considered essential to guarantee the reliability of neuroimaging studies. It is particularly important for fetal brain MRI, where acquisitions and image processing techniques are less standardized than in adult imaging. In this work, we focus on automated quality control of super-resolution reconstruction (SRR) volumes of fetal brain MRI, an important processing step where multiple stacks of thick 2D slices are registered together and combined to build a single, isotropic and artifact-free T2 weighted volume. We propose , a machine-learning method that extracts more than 100 image quality metrics to predict image quality scores using a random forest model. This approach is well suited to a problem that is high dimensional, with highly heterogeneous data and small datasets. We validate in an out-of-domain (OOD) setting and report high performance (ROC AUC = 0.89), even when faced with data from an unknown site or SRR method. We also investigate failure cases and show that they occur in of the images due to ambiguous configurations for which the rating from the expert is arguable. These results are encouraging and illustrate how a non deep learning-based method like is well suited to this multifaceted problem. Our tool, along with all the code used to generate, train and evaluate the model are available at https://github.com/Medical-Image-Analysis-Laboratory/fetmrqc_sr
@@
Although the quality of SRR volumes greatly influences subsequent processing steps like automatic tissue segmentation and surface extraction, only a few methods have been proposed, and their validation was limited. Closest to this work, the study of Largent et al. [18] used multi-instance deep learning methods for automated quality control on 3D volumes reconstructed using only the approach from Jiang et al. [10]. The model trained in this work is not publicly available and the validation was restricted to one acquisition site. Other works have taken a more descriptive approach to the quality of SRR volumes. The work of Rubert et al. [19] focused on one popular SRR method called NiftyMIC [14], and proposed to predict the quality of the reconstructed volume based on the percentage of slices from the input stacks that NiftyMIC uses for reconstructing that volume. The work of Sanchez et al. [17] proposed a protocol for fetal brain MRI quality assessment based on the evaluation of several image and artifact properties, alongside an overall image quality score [20]. However, their sample size remained quite small and they did not perform an automated quality assessment. Finally, while a few automated QC methods have been proposed for low-resolution fetal brain MRI [21, 22, 23], they do not readily apply to SRR volumes, since the artifacts between raw T2-volumes and SRR volumes are very different.
In this work, we propose to tackle the challenge of automated QC for multi-SRR, multi-site data. Our contributions are as follows: (i) We collect a unique dataset consisting of 673 manual quality ratings combining multiple acquisition sites and multiple SRR methods (ii) We build upon FetMRQC [23] to propose a robust approach based on random forest. We show that this non deep-learning approach, when combined with appropriate data pre-processing, provides a high performance, while requiring only a modest amount of training data. (iii) We evaluate the out-of-domain robustness of our method, focusing on unseen sites and SRR methods. (iv) We study the failure modes of our model and show that 45% of the failures are due to ambiguous ratings, and that nearly all false negatives (wrongly flagged as bad quality) have some level of artifact that, although it is not disturbing to a human rater, is picked up by . The code is publicly available on GitHub111https://github.com/Medical-Image-Analysis-Laboratory/fetmrqc_sr. We believe that this method will serve as the first open source, easy-to-use baseline for future works in the field.
@@ 2.1 Demographics
Data. Fetal brain MRI examinations were retrospectively collected from ongoing research studies at three hospitals: CHUV (n=141), BCNatal (n=180) and KISPI (n=37). The study received ethical approval from each center’s institutional review board. These 358 subjects were acquired on different Siemens scanners (Erlangen, Germany) at 1.5T or 3T for CHUV and BCNatal, and on different GE Healthcare (Chicago, USA) scanners for KISPI. As the aim of quality control is to be applicable to a wide range of scenarios, each subject was reconstructed using a randomly assigned set of parameters: a randomized SRR method, target resolution () and selection of input stacks. This resulted in 673 reconstructed volumes using four widely-used SRR methods: SVRTK [11, 24] (n=238), NeSVoR [15] (n=213), NiftyMIC [14] (n=202), MIALSRTK [13] (n=20). Additional details regarding the number of stacks used per reconstruction, echo and repetition times are provided in Table S1.
Manual quality ratings. After the SR reconstructions were completed, a manual quality assessment was performed following the protocol of Bach Cuadra et al. [20]. The protocol was slightly adapted to focus on categories that were found to have an inter-rater reliability of , leading to the categories illustrated on the right of Figure 1. Global quality rating was obtained as a continuous variable between 0 and 4, and the exclusion threshold was set to 1. Two experts, M.B.C. (20 years of experience in fetal brain MR image analysis) and T.S. (4 years of experience) performed manual quality annotations on all the reconstructions (323 for M.B.C. and 448 for T.S., with 98 cases rated twice). The annotation interface allowed to streamline the process, and ratings took a total of 11h40 (7h10 for M.B.C., 4h30 for T.S.), with a median time of 47 seconds per case. Details on the distribution of SRR quality for the three main methods are provided in Figure 3.
@@ 2.2 Automated quality control
Image quality metrics. IQMs are features capturing complementary statistics on a given volume. Well-designed IQMs correlate with various image artifacts, enabling the training of a predictive QC model. However, IQMs that are commonly used in adult MRI are not readily applicable to fetal SRR volumes, as they can rely for instance on air as a background [5], which is inappropriate in utero as the fetus is surrounded by maternal tissues. This prompted us to adapt the IQMs of FetMRQC [23], designed for stacks of 2D slices of T2w fetal brain MRI, to the SRR setting. These IQMs are based on image intensity (e.g. mean intensity, SSIM of neighbouring slices, etc.), mask statistics (brain mask centroid coordinates) or segmentation-related statistics (tissue-wise statistics like volume, or contrast between tissue intensities). In addition to these metrics we incorporate new topological IQMs based on Betti numbers and Euler characteristics [25], as this has been shown to be an important metric to quantify the quality of a fetal brain MRI segmentation [26, 27]. Since several IQMs required a brain mask and tissue segmentation, we used the state-of-the-art BOUNTI [28] to compute these maps. In total, 106 IQMs were computed (28 from FetMRQC, 54 from segmentation and 24 from topology). A detailed description of the IQMs used is provided in Table S2.
@@
Predictive model. is based on a random forest, and trained in a binary classification setting, where it predicts whether a data sample has quality below the threshold of , and should be excluded. In addition to it, we considered two variants of . The first one was trained by re-weighted samples according to their quality score. Given a sample-label pair , the continuous quality score ranging from 0 (exclude) to 4 (excellent) was mapped to a 0-1 probability of being of good quality by using a sigmoid function centered at the decision boundary (quality score=1):
The sample then received a weight , based on its label, computed as the purity ( 1 - the binary entropy ) of label :
We then have , as these labels correspond to very unambiguous examples, and (minimal purity), as a sample at the decision boundary is inherently uncertain. It is referred to as reweighting, and shown in Figure 2. The second one, called Predict_artifacts, made use of the available manual ratings of the type of artifacts present: random forest regressors were trained to predict each specific artifact using the IQMs (cf. Figure 1 right), and then these predictions were used as additional IQMs for the prediction.
Implementation. was implemented using Python 3.9.0, scikitlearn 1.1.3, Gudhi 3.11.0 for topology computation and called on a docker wrapper of BOUNTI222fetalsvrtk/segmentation:general_auto_amd. It builds on the code-base of FetMRQC [23]333https://github.com/Medical-Image-Analysis-Laboratory/fetmrqc.
@@ 2.3 Experiment design
2.3.1 Evaluation setting.
Previous studies have shown that domain shift is a major obstacle to the generalization of ML-based methods in medical imaging [29], motivating us to evaluate our models in four different experiments. In-domain (ID), where the data were split by subject across centers in 10-fold cross-validation (CV); leave-one-site-out CV, where all the data from a single site were used for testing; leave-one-SRR-out CV, where all the data from a given SRR were used for testing; leave-one-site/SRR-out, where all the data from a given site and a given SRR method were left out during training, and the model was tested on the combined data of the left-out site and SRR. This simulated a complex domain shift, where our method was tested on a new site, with data reconstructed using an unseen SRR method. Each experiment was repeated 10 times to account for random variations in the model training and training data sampling. All experiments used the same 673 samples but relied on different data splits. No pure testing evaluation was included in this paper due to space constraints.
@@ 3 Results
We first carried out an in-domain ablation (subjects split using 5-fold CV) study on the various preprocessing options and variants of . We also considered various thresholds for exclusion from the model’s predicted probabilities, and investigated how changing impacted BA, sensitivity and specificity. Results are presented in Table 1, where we see that using segmentation and IQMs pre-processing reaches the best performance with AUC=0.93 and BA=0.85. Choosing increases the specificity of the model, and we chose to keep it at that level for the rest of the paper. The addition of topological metrics improved the AUC and BA by 3% and 4% respectively. Finally, when looking at the oracle, which uses manual ratings of artifacts as the basis for prediction, we see that our models come close to the predictive ability of the oracle, but fall short by a margin of approximately 7-8%. We also experimented with the site-SRR domain shift removal using ComBaT [30], robust scaling, alternative classifiers like logistic regression, and nested cross-validation but these approaches did not improve the performance and are not reported. An analysis of the feature importances of in a 10-fold CV setting is presented on Figure S5.
@@
In this experiment, we evaluate our model on three OOD settings: leave-one-site-out, leave-one-SRR-out and leave-one-site/SRR-out CV. Our results, in Table 2, show an increasing drop in performance across settings: it is less challenging to generalize to an unseen site than to an unseen SRR, and the most challenging setting is the double shift site+SRR. Although the AUC and BA do not vary much for with robust preprocessing, there is a large drop in sensitivity when evaluating the model on a different site (-7%) and SRR (-11%) and a drop in both sensitivity and specificity in the site+SRR case (resp. -9%; -13%). The Site+SRR case also shows a larger variance in sensitivity and specificity of models, suggesting that the model might exhibit some degree of unreliability on this challenging domain shift. Nevertheless, generally retains the best average performance.
@@
Out of 673 cases, the prediction was consistently false for 95 cases (14%) across the repeated experiments. Figure 4 breaks down the distribution of the errors as a function of the manual quality rating. Out of the 95 failed cases, 23 correspond to cases with a strong bias field, and 22 to a low tissue contrast (Figure 4D). Both were challenging for ’s IQMs to capture and accurately classify. This is one limitation of feature engineering approaches that cannot capture all artifacts encountered. This is further illustrated by the cases with streaks and cut brain (Figure 4A and B), that are confidently predicted as good quality, while being excluded by raters. Nevertheless, out of 95 failed cases, 42 had reference quality scores close to the decision boundary and we consider them as ambiguous cases, where a different rater could have assigned a different label (cf. Figure 4C). These cases might be interpreted as limitations from the ground truth data as well as a challenging cases for . Lastly, systematic failure cases also included images of obviously pathological cases, for which the automated segmentation method failed, leading in turn to inaccurate IQMs and mis-classification by the model.
@@ 4 Discussion
While not reported, we also experimented with CNN-based models[31], which we found to be challenging this setting: training was very unstable, and we could not achieve stronger performance than . We believe that DL-based models could be developed for this application, but found it to be not as easy as we would have expected and leave it for future work. Finally, we believe that including additional IQMs based on radiomics features [32] could also be an interesting avenue in the future.
@@ 4.0.1 Acknowledgements
This work was funded by Era-net NEURON MULTIFACT project (TS: Swiss National Science Foundation grant 31NE30_203977; AM, GA: French National Research Agency, Grant ANR-21-NEU2-0005; EE: Instituto de Salud Carlos III (ISCIII) grant AC21_2/00016 , GMJ: Ministry of Science, Innovation and Universities: MCIN/AEI/10.13039/501100011033/), the SNSF project 215641, and the SulcalGRIDS Project, (GA: French National Research Agency Grant ANR-19-CE45-0014). AJ is supported by the Prof. Max Cloetta Foundation, EMDO Foundation and Vontobel Foundation. This research was also supported by grants from NVIDIA through an Academic Grant Program and utilized the provided NVIDIA RTX6000 ADA GPUs. TS, VZ and MBC acknowledge access to the facilities and expertise of the CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by CHUV, UNIL, EPFL, UNIGE and HUG.
@@
Supplementary material
| Site | Scanner | SRR | TE [s] | TR [s] | Thickness [mm] | ||||
BCNatal |
Aera (1.5T) | NeSVoR | 35 | 11.17±4.65 | 0.10±0.03 | 1.27±0.25 | 2.87±0.44 | ||
| NiftyMIC | 20 | 8.10±3.99 | 0.10±0.02 | 1.18±0.24 | 2.98±0.41 | ||||
| SVRTK | 20 | 9.25±4.18 | 0.10±0.02 | 1.22±0.24 | 2.98±0.44 | ||||
| TrioTim (3T) | NeSVoR | 55 | 4.73±1.96 | 0.14±0.00 | 1.27±0.43 | 3.83±0.66 | |||
| NiftyMIC | 53 | 4.28±1.50 | 0.14±0.00 | 1.39±0.73 | 3.76±0.57 | ||||
| SVRTK | 62 | 5.34±3.21 | 0.14±0.00 | 1.38±0.74 | 3.74±0.64 | ||||
| Other | NeSVoR | 7 | 4.00±0.82 | 0.08±0.01 | 1.42±0.20 | 3.14±0.24 | |||
| NiftyMIC | 5 | 4.20±1.64 | 0.08±0.00 | 1.55±0.03 | 3.00±0.00 | ||||
| SVRTK | 6 | 5.00±1.41 | 0.08±0.00 | 1.51±0.10 | 3.00±0.00 | ||||
CHUV |
Aera (1.5T) | MIALSRTK | 9 | 6.78±2.05 | 0.13±0.11 | 1.37±0.51 | 2.86±0.43 | ||
| NeSVoR | 47 | 7.68±3.67 | 0.10±0.07 | 1.26±0.31 | 2.99±0.29 | ||||
| NiftyMIC | 62 | 6.61±2.77 | 0.10±0.06 | 1.25±0.27 | 2.99±0.30 | ||||
| SVRTK | 60 | 8.12±3.87 | 0.10±0.06 | 1.25±0.28 | 3.03±0.35 | ||||
MIALSRTK | 8 | 6.50±1.41 | 0.09±0.00 | 1.20±0.00 | 3.00±0.00 | |||
| NeSVoR | 37 | 7.73±3.82 | 0.10±0.06 | 1.24±0.25 | 2.98±0.22 | ||||
| NiftyMIC | 32 | 7.34±3.83 | 0.09±0.00 | 1.20±0.00 | 3.00±0.02 | ||||
| SVRTK | 46 | 6.61±2.44 | 0.11±0.09 | 1.30±0.38 | 2.93±0.33 | ||||
| Other | MIALSRTK | 3 | 7.00±1.73 | 0.09±0.01 | 1.07±0.06 | 3.33±0.58 | |||
| NeSVoR | 5 | 7.00±2.35 | 0.10±0.00 | 1.10±0.00 | 3.04±0.09 | ||||
| NiftyMIC | 11 | 6.73±1.85 | 0.10±0.01 | 1.09±0.03 | 3.09±0.30 | ||||
| SVRTK | 8 | 8.62±4.27 | 0.10±0.00 | 1.10±0.00 | 3.00±0.00 | ||||
KISPI |
NeSVoR | 2 | 4.50±2.12 | 0.12±0.00 | 3.00±0.00 | 3.00±0.00 | ||
| NiftyMIC | 5 | 8.80±4.15 | 0.11±0.01 | 2.32±0.92 | 2.80±0.27 | ||||
| SVRTK | 5 | 14.80±3.27 | 0.11±0.01 | 2.45±0.51 | 4.60±1.52 | ||||
NeSVoR | 6 | 7.67±1.97 | 0.12±0.00 | 3.00±0.00 | 3.00±0.00 | |||
| NiftyMIC | 2 | 8.50±2.12 | 0.12±0.00 | 3.00±0.00 | 3.00±0.00 | ||||
| SVRTK | 5 | 7.20±2.28 | 0.11±0.01 | 2.68±0.72 | 2.90±0.22 | ||||
NeSVoR | 13 | 13.38±6.38 | 0.12±0.01 | 2.49±1.08 | 3.08±0.61 | |||
| NiftyMIC | 9 | 13.33±4.09 | 0.12±0.01 | 2.04±0.94 | 3.17±0.71 | ||||
| SVRTK | 24 | 15.62±7.27 | 0.12±0.01 | 2.46±0.92 | 2.98±0.48 | ||||
NeSVoR | 6 | 13.83±4.36 | 0.11±0.01 | 2.55±0.50 | 4.00±1.10 | |||
| NiftyMIC | 3 | 9.67±2.52 | 0.11±0.01 | 2.15±0.81 | 3.50±1.32 | ||||
| SVRTK | 2 | 15.50±7.78 | 0.11±0.02 | 2.52±0.67 | 4.00±1.41 |
| Intensity-based IQMs | ||
| IQMs | Count | Description |
| rank_error [12] | 2 | Measure the compressibility of the image using a low-rank approximation. Two variants: absolute and relative rank error [12]. |
| slice-wise difference [12, 11, 14] | 14 | Metrics for outlier rejection: normalized MAE, mutual information, cross-correlation, RMSE, PSNR, SSIM, joint entropy. Two variants to each metric: Mean/Median of differences across slices computed within 3 slice window (7 metrics 2 variants) |
| entropy [5] | 1 | Entropy across the brain volume |
| sstats [5] | 7 | Mean, median, std, percentiles (5%, 95%), coefficient of variation, kurtosis on brain ROI |
| filter_image [23] | 2 | Sharpness estimation using Laplace and Sobel filters |
| Masked-based IQMs | ||
| mask_volume [23] | 1 | Volume of the brain mask |
| centroid [35] | 1 | Location of the brain mask centroid |
| Segmentation-based IQMs | ||
| sstats [5] | 40 | Stats per region (WM, GM, CSF, BS, CBM): mean, median, 5th/95th percentile, kurtosis, std, MAD, voxel count (5 regions 8 stats) |
| volume [5] | 5 | Volume of WM, GM, CSF, BS, CBM |
| SNR [36] | 6 | Signal-to-noise ratio per region and globally |
| CNR [37] | 1 | Contrast-to-noise ratio between GM and WM |
| CJV [38] | 1 | Coefficient of joint variation of GM and WM |
| WM2MAX [5] | 1 | Ratio of white matter to maximum intensity |
| Topology-based IQMs (WM, GM, CSF, BS, CBM and whole brain) | ||
| Betti numbers BN0, BN1, BN2 [25] | 18 | Betti numbers describing the number of connected components (BN0), the number of loops or holes (BN1) and the number of cavities (BN2) for five brain regions (WM, GM, CSF, BS, CBM) and entire brain (3 Betti numbers 6 regions) |
| Euler Characteristic [25] | 6 | Euler characteristic (= BN0 - BN1 + BN2) for each region and entire brain |
![[Uncaptioned image]](x4.png)